I need in slope intercept form please
1) the line passing through (-4,6) and parallel to the line y=3x+8
2) the line passing through (1,2) and perpendicular to the line y=4x+3

Answer by Alan3354(69443)

I need in slope intercept form please
1) the line passing through (-4,6) and parallel to the line y=3x+8
The given lines are in slope-intercept form, y = mx + b
m = slope
b = y-intercept
---------------
All parallel lines have the same slope
The slope of the eqn will be 3
Use the slope and and the point to find b
---
6 = 3*(-4) + b
6 = -12 + b
b = 18
--> y = 3x + 18
==========================
2) the line passing through (1,2) and perpendicular to the line y=4x+3
The slope of the given line m = 4
Lines perpendicular have a slope that the negative inverse, so m = -1/4
The rest is the same as Problem #1